Circular No. 8613

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/1993 W1 = 2005 T1 (MUELLER) E. J. Christensen has reported the recovery of comet P/1993 W1 (=

1993s = 1994 XXV) on CCD images taken with the 1.5-m reflector in the course of the Mt. Lemmon Survey, describing the object as having a condensed 8′′ coma and a 20′′ curved tail in p.a. 250^{o}–280^{o}. The recovery was confirmed at the Minor Planet Center via LONEOS incidental astrometry (0.59-m Schmidt telescope) by M. E. Van Ness, and earlier observations were subsequently also reported by F. Fratev (Zvezdno Obshtestvo Observatory, Plana, 0.25-m f/4.7 reflector). The indicated correction to the prediction on MPC 54168 is ΔT = −0.8 day.

2005 UT α_{2000} δ_{2000} Mag. Observer
Oct. 6.93268 2^{h}39^{m}18.^{s}35 −3^{o}55′12′′.2 18.4 Fratev
6.94519 2 39 18.06 −3 55 14.5 18.4  _{”}
6.95448 2 39 17.90 −3 55 14.5 18.4  _{”}
6.96144 2 39 17.56 −3 55 16.2 18.4  _{”}
7.32842 2 39 08.90 −3 56 21.4 18.5 Van Ness
7.35758 2 39 08.13 −3 56 26.4  _{”}
7.41793 2 39 06.68 −3 56 36.5  _{”}
7.43542 2 39 06.30 −3 56 39.4 18.0 Christensen
7.44722 2 39 05.99 −3 56 41.7  _{”}
7.44923 2 39 06.01 −3 56 41.7  _{”}

The following orbital elements by B. G. Marsden, Smithsonian Astrophysical Observatory, are from 83 observations spanning 1992–2005 (mean residual 0′′.8):

Epoch = 1994 Sept. 5.0 TT

T = 1994 Sept.12.1603 TT  ω = 30.^{o}0222
e = 0.260528  Ω = 100.6625 }2000.0
q = 4.249738 AU  i = 16.4655

a = 5.746990 AU n^{o} = 0.0715391 P = 13.78 years Epoch = 2008 May 14.0 TT

T = 2008 May 18.5222 TT  ω = 29.^{o}8348
e = 0.261149  Ω = 100.5697 }2000.0
q = 4.214478 AU  i = 16.4959

a = 5.704097 AU n^{o} = 0.0723475 P = 13.62 years

2005 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green